OpenStudy (anonymous):

What is the factored form of the following expressions? x2 – 10xy + 24y2

OpenStudy (anonymous):

-10=-4-6 24=(-4)(-6) x2 – 10xy + 24y2 =(x-4y)(x-6y)

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x^2 - 4xy - 6x + 24y^2 }\) \(\Large \color{MidnightBlue}{\Rightarrow x(x - 4y) - 6(x - 4y) }\) \(\Large \color{MidnightBlue}{\Rightarrow ( x - 6)(x - 4y) }\)

Parth (parthkohli):

@CoCoTsoi you might wanna check it again

OpenStudy (anonymous):

@ParthKohli thank you!

OpenStudy (anonymous):

What happened to my answer? :D For @parth 2nd step, -6x-24y^2 is not equal to -6(x-4y)

OpenStudy (anonymous):

@ParthKohli

OpenStudy (ajprincess):

ur answer is right @CoCoTsoi

OpenStudy (anonymous):

thanks, @ajprincess

OpenStudy (ajprincess):

yw.